OpenStudy (anonymous):

The school cafeteria sells two kinds of wraps: vegetarian and chicken. The vegetarian wrap cost $1 and the chicken wrap cost$1.80. Today they made $98.80 from the 70 wraps sold. How many of the wraps sold were vegetarian?

OpenStudy (anonymous):

let: x = the number of chicken wrap y = vegi wrap 70 = x + y --------> 1st equation 98.8 = x(1.8) + y(1) --------> 2nd equation from 1st equation: x = 70 - y ---------> substitute to the 2nd equation. 98.8 = (70-y)(1.8) + y(1) -----> solve for y. 98.8 = 126 - 1.8y + y 98.8 = 126 - 0.8y 0.8y = 126 - 98.8 0.8y = 27.2 y = 27.2/.8 y = 34 ------> (34 vegetarian wraps where sold)

OpenStudy (anonymous):

something was telling me that it was 36

OpenStudy (anonymous):

check it. :))

OpenStudy (anonymous):

thank u 4 your help always!!!

OpenStudy (anonymous):

yeah kumarsajan is right. the one who told you that the answer is 36, he meant the chicken warps. 34+ 36 = 70 :)